OpenStudy (anonymous):

Please help, I'll give you a medal<3 Below are the results of tossing a number cube 8 times. Find the experimental probability of tossing a an even number. 3,4,1,2,5,1,6,5 A. 1/4 B.3/8 C.1/2 D.5/8

jimthompson5910 (jim_thompson5910):

how many even numbers are in that list?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Is it B?

jimthompson5910 (jim_thompson5910):

yep there are 3 numbers you want (4,2,6) out of 8 total

OpenStudy (anonymous):

Thank you(:

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Help with another?

jimthompson5910 (jim_thompson5910):

sure, one more

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (anonymous):

A bag contains tiles with the letters C-O-M-B-A-N-A-T-I-O-N-S. Lee chooses a tile without looking and doesn't replace it.

OpenStudy (anonymous):

He chooses a second tile without looking. What is the probability that he will choose the letter O both times?

OpenStudy (anonymous):

A. 1/132 B. 1/72 C.1/66 D.1/23

jimthompson5910 (jim_thompson5910):

What is the probability he chooses an O the first time?

OpenStudy (anonymous):

1/12?

jimthompson5910 (jim_thompson5910):

not 1/12 but close

OpenStudy (anonymous):

Um... 2/12

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

once you pick an O out, you are left with 2-1 = 1 O tile left this is out of 12-1 = 11 tiles

jimthompson5910 (jim_thompson5910):

You then multiply the two probabilities (2/12)*(1/11) = ???

OpenStudy (anonymous):

It's 2/23?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

multiply, not add

jimthompson5910 (jim_thompson5910):

reduce as much as possible

OpenStudy (anonymous):

oh opps, lmao. It's 2/132

OpenStudy (anonymous):

So it would equal 1/66.

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

THANK YOU!!!!

jimthompson5910 (jim_thompson5910):

np